I am not sure I got this right:
Divide and simplify.
(7x-56)/55 divided by (x-8)/33x
I got two different answers, which can't be right.

Found 3 solutions by sudhanshu_kmr, Alan3354, MathLover1:
Answer by sudhanshu_kmr(1152)


(7x-56)/55 divided by (x-8)/33x
= (7x-56)/55 * 33x / (x-8)
= 7*33x /55
= 21x/5

Answer by Alan3354(69443)

(7x-56)/55 divided by (x-8)/33x
= (7x-56)/55 * 33x/(x-8)
= 7(x-8)/55 * 33x/(x-8)
= 7*33x/55
= 21x/5




Answer by MathLover1(20849)